A company with a share price of $49 has shares outstanding of 5 million and net income of $20 million. The company conducts a leverage recapitalization to alter its capital structure by borrowing $3 million and using the proceeds to buy back shares. The pre-tax cost of debt is 6% and the company's tax rate is 29%.

What will be the new earnings per share after the leverage recapitalization?

Enter your answer rounded to two decimals.
Answer #1

first of all lets find out the profit before tax = 20 Million/(1-29%)

= 28169014.08451

Interest on Loan = 3000000*6% = 180000

Profit Before tax = 28169014.08451-180000 = 27989014.08

Profit After Tax = 27989014.08*(1-0.29) = 19872200

EPS =19872200/5000000 = $ 3.97 per share.